OpenStudy (anonymous):

For the game she's playing, Liz must draw a card from a a deck of 26 cards,one with each letter of the alphabet on it,and roll a six-sided die.What is the probability that Liz will roll an old numbetr and draw a letter in her name? page 144

OpenStudy (anonymous):

\[\frac{3}{26}\times \frac{1}{2}=\frac{3}{52}\]